OpenStudy (anonymous):

Simplify: cube root of 5 over fourth root of 5

OpenStudy (anonymous):

@Euler271 may I have some help

OpenStudy (anonymous):

@kirbykirby

OpenStudy (anonymous):

\[\frac{ x^a }{ x^b } = x^{a-b}\]

OpenStudy (anonymous):

\[\sqrt{x} = x^{1/2}\]\[\sqrt[3]{x} = x^{1/3}\]\[\sqrt[n]{x} = x^{1/n}\]

OpenStudy (anonymous):

\[\frac{ \sqrt[3]{5} }{ \sqrt[4]{5} } = \frac{ 5^{1/3} }{ 5^{1/4} } = 5^{1/3 - 1/4} = 5^{4/12 - 3/12} = 5^{1/12} = \sqrt[12]{5}\]

OpenStudy (anonymous):

I'm sorry, would you mind drawing it out, it didn't translate the information you gave me. :(

OpenStudy (anonymous):

it's kind of long to write out lol. your browser must be the problem. you should try closing and re-opening it (copy link first to paste it) or using a different one :P

OpenStudy (anonymous):

sometimes works even to just refresh the page

OpenStudy (anonymous):

Alright, hold on one second.

OpenStudy (anonymous):

Huh, it didn't work.

OpenStudy (anonymous):

Well, could you explain in like a couple of words how you got 5 1/12?

OpenStudy (anonymous):

I just want to know how you did it.

OpenStudy (anonymous):

OH MY GOSH. You found the common denominators of 3 and 4 and THEN you subtracted, duh! Gosh, I need to practice more. Thank you! XD

OpenStudy (anonymous):

happy to help :) i forgot about how convenient it was to upload pictures >.<

OpenStudy (kirbykirby):

I'm having the same problem too where LaTeX won't show up anymore. I'm using the same browser as I always have... o_O

OpenStudy (anonymous):

i can't use draw >.<

OpenStudy (kirbykirby):

actually Draw isn't working for me either also :o